The pH of a 0.160 M solution of a weak base is 9.21. What is the Kb of the base?


1
Expert's answer
2021-12-13T04:49:02-0500

If pH = 9.21


pOH = 14.0 - 9.21 = 4.79


Then [OH-] = 10^-4.79


[OH-] = 6.31*10^-5


This is low enough to allow the use of the short cut method


If the base is MOH


Then Kb = [M+] [OH-] / [MOH]


[M+] = [OH-] and [MOH] = 0.1 M


Kb = (6.31*10^-5)² / 0.1


Kb = 3.98*10^-9 / 0.1


Kb = 3.98*10^-8